Assume that a sample is used to estimate a population proportion p. Find the 99% confidence interval for a sample of size 308 with 123 successes. Enter your answer as a tri-linear inequality using decimals (not percents) accurate to three decimal places.

< p <



Answer should be obtained without any preliminary rounding. However, the critical value may be rounded to 3 decimal places.
Answer #1

Confidence interval for Population Proportion is given as below:

Confidence Interval = P ± Z* sqrt(P*(1 – P)/n)

Where, P is the sample proportion, Z is critical value, and n is sample size.

We are given

x = 123

n = 308

P = x/n = 123/308 = 0.399350649

Confidence level = 99%

Critical Z value = 2.5758

(by using z-table)

Confidence Interval = P ± Z* sqrt(P*(1 – P)/n)

Confidence Interval = 0.399350649 ± 2.5758* sqrt(0.399350649*(1 – 0.399350649)/308)

Confidence Interval = 0.399350649 ± 2.5758* 0.0279

Confidence Interval = 0.399350649 ± 0.0719

Lower limit = 0.399350649 - 0.0719 = 0.3275

Upper limit = 0.399350649 + 0.0719 = 0.4712

Confidence interval = (0.3275, 0.4712)

0.3275 < p < 0.4712